# One Step is Enough: Multi-Agent Reinforcement Learning based on One-Step Policy Optimization for Order Dispatch on Ride-Sharing Platforms

### v1

@@ Introduction
If we ignore the residual term, we obtain $\hat{\text{A}}_{\pi}(s_{t},u_{t})\approx\frac{1}{\sigma}(\text{Q}_{\pi}(s_{t},u_{t})-\text{V}_{\pi}(s_{t}))$. Compared to the advantage in PPO (Eq. 2), there is only one additional factor $\frac{1}{\sigma}$. According to (Kurin et al. 2022), this can be viewed as normalization for the reward function, contributing to reduced deviation and more stable training. Furthermore, based on the theory of policy gradients (Sutton et al. 1999), adding any baseline or positive factor does not influence the policy direction, regardless of the action. As a result, we consider the advantage of GRPO (Eq. 3) to be a valid replacement for the advantage in PPO (Eq. 2).
One Step Policy Optimization (OSPO)
We note that in the ride-sharability platform, all vehicles have identical capacity and are assigned to customers purely based on spatial proximity, without any additional attributes or preferences influencing the dispatch decisions. As a result, in the long run, the cumulative contribution or average earnings of each driver per unit time must converge to the same value. This is because the inherent randomness in order arrivals and assignments is experienced equally by all drivers over time. Therefore, it is reasonable to expect that different agents will exhibit similar value functions at each time step, regardless of their individual states. In this case, we have the following property:
$\text{V}_{\pi}(s_{1,t})+\epsilon_{1,t}=\text{V}_{\pi}(s_{2,t})+\epsilon_{2,t}=\cdots$ (8)
$=\text{V}_{\pi}(s_{n,t})+\epsilon_{n,t}=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[\text{V}_{\pi}(s_{i,t})]\ ,$
where $\epsilon_{i,t}$ is a small residual term, with $\mathbb{E}_{i\in\mathcal{I}}[\epsilon_{i,t}]=0$. As a result, when we ignore the residual terms, the different agents do not differ significantly from the group in GRPO, making it feasible in our scenario. Based on the property in Eq. 8, we can recalculate the advantage function as follows:
@@
If we perform a similar operation to GRPO, using the average reward to replace the expectation, applying reward normalization (Kurin et al. 2022), and ignoring the residual term, we arrive at the advantage function:
@@
This indicates that we can use only the one-step reward for policy optimization, which we term One Step Policy Optimization (OSPO). Even though the format may seem counter-intuitive, we provide an intuitive explanation. The validity of Eq. 10 is based on Eq. 8: the V-values remain similar among agents at each step due to the homogeneity in the ride-sharing system. For any arbitrary agent, regardless of its actions in the current step, it can achieve a similar V-value as others in the next step. Therefore, under the independent assumption, the goal of each agent is to maximize its single-step reward. In Eq. 8, removing the baseline and factor leaves only the single-step reward $\text{r}(s_{i,t},u_{i,t})$, which aligns with our intuition. Compared to GRPO, our OSPO further improves computational efficiency. Moreover, our method demonstrates greater robustness since it is not influenced by the discount factor.
@@
$\bar{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}\ .$ (11)
@@
This approach not only makes $\mu_{t}$ closer to $\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]$ in Eq. 9, but also further improves computational efficiency, achieving true one-step computation. Second, during training, to better align with the property described in Eq. 8, we propose a penalty term $\Delta\delta_{t}$, where $\delta_{t}$ represents the standard deviation of the current cumulative reward for each agent, and $\Delta\delta_{t}=\delta_{t+1}-\delta_{t}$. This penalty is designed to reduce the utility gap (reflecting the V-value) between agents. Consequently, the final advantage function is expressed as:
$\tilde{\text{A}}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}-\alpha\Delta\delta_{t}\ ,$ (12)
@@
where $\alpha$ is a hyper-parameter. (Note that when Eq. 8 is satisfied, $\Delta\delta_{t}$ is close to 0. We also include the term $-\alpha\Delta\delta_{t}$ in the advantage function of GRPO.)
Experiment
Experiment Setup
@@
To validate our method, we conduct a series of experiments using a real-world ride-hailing dataset from Manhattan, New York City (Taxi and Commission 2024). The data collected from 19:00 to 19:30 on July 17, 2024, is used as the training set, which includes 3,726 valid orders. The data from other periods of that day is used as the testing set.
@@
For the simulation, we set the total number of drivers to 1,000, with each car having a capacity of 3 and a default speed of 60 km/h. The optimal routing is achieved using the OSRM simulator (Luxen and Vetter 2011). The experiments are conducted using the PyTorch framework (Paszke et al. 2019) on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card. Due to device limitations, we set each episode length to 30 minutes, with each step representing 1 minute, resulting in a running time of approximately 40 to 120 seconds for different methods per episode. We maximize the training episodes at 1,000 and evaluate model performance every 10 episodes.
@@
To further illustrate the efficiency of our method, we choose a simple four-layer MLP as the policy network, with 128 units in each hidden layer and LeakyReLU (Maas et al. 2013) as the activation function. Considering the variable action space (since the number of orders changes each time), we set the network input as a single driver-order pair, ignoring the relationships between orders. Finally, a softmax layer is used to normalize the output probabilities.
@@
During training, our models utilized approximately 3.8GB (OSPO) and 5.2GB (GRPO) of GPU memory, with a batch size of 256. For exploration, we add gradually decreasing random noise, employing the same Binary Symmetric Channel (BSC) noise as in (Hu, Feng, and Li 2025). For optimization, we use the Adam optimizer (Kingma 2014) with an initial learning rate of $10^{-4}$ and a decay rate of 0.99.
@@
In the following, we first compare our GRPO and OSPO ride-sharing methods with previously popular methods of different types, including DTDE, CTDE, and CTCE. Next, we conduct a series of ablation studies to detail the effects of different modules in GRPO and OSPO, comparing them with similar policy gradient-based methods, including Independent PPO (IPPO) (De Witt et al. 2020), Multi-Agent PPO (MAPPO) (Yu et al. 2022), and Independent PG (IPG) (Sutton et al. 1999).
Comparative Experiment
Figure 2: Training Process: Each method is trained three times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$. The shaded area represents the range of fluctuations, while the solid line indicates the average value. (For delivery time and detour time, only completed orders are counted, as these metrics are uncertain for unfinished orders.)
@@
Table 1: Model Performance on Testing Sets: Bold entries represent the best results.
Method Configuration Order Amount in Testing Sets Method Network GPU (GB) 2,850 3,114 3,577 3,910 DeepPool MLP (20K) 3.97 10,690.95±86.93 12,368.83±117.75 12,569.92±77.15 12,990.07±25.07 BMG-Q GAT (117K) 4.28 10,982.45±187.60 12,637.12±116.89 13,064.66±142.65 13,600.33±92.30 HIVES GRU (16M) 6.01 11,044.05±8.30 12,787.39±40.97 12,576.75±46.25 12,097.87±72.32 Enders et al. MLP (118K) 8.19 10,036.89±88.64 12,160.56±39.53 12,296.22±188.86 12,128.95±123.22 CEVD MLP (23K) 21.45 10,842.98±31.36 12,702.10±37.39 13,263.08±136.67 13,609.91±17.34 GRPO MLP (20K) 5.17 11,372.81±31.73 13,015.88±78.18 13,644.95±116.94 12,964.86±99.84 OSPO MLP (20K) 3.82 11,467.87±5.39 13,046.27±15.29 13,771.56±81.43 13,429.67±69.79
@@
To illustrate the efficiency of our methods, we compare them with several representative ride-sharing approaches from various categories. A more detailed overview of previous work is provided in the appendix.
@@
  • DeepPool (Al-Abbasi, Ghosh, and Aggarwal 2019) (DTDE): DeepPool is one of the earliest MARL-based ride-sharing works, based on IDDQN (Tan 1993; Van Hasselt, Guez, and Silver 2016) and using a CNN as the Q-network. To align it with our experimental scenario, we replace the CNN with an MLP similar to our methods (without the softmax).
@@
  • BMG-Q (Hu, Feng, and Li 2025) (DTDE): To achieve better cooperation among agents, BMG-Q is also based on IDDQN but utilizes Graph Attention Network (GAT) (Velickovic et al. 2017) to capture the relationships among neighboring agents.
@@
  • HIVES (Hao and Varakantham 2022) (CTDE): Based on QMIX (Rashid et al. 2020), HIVES tries addressing the curse of dimensionality problem by proposing a novel hierarchical mixing structure.
@@
  • Enders et al. (Enders et al. 2023) (CTDE): Based on Multi-Agent Soft Actor Critc (MASAC) (Haarnoja et al. 2018), Enders et al. propose allowing agents to choose whether to accept assigned orders, which helps the system serve more valuable orders instead of treating all orders equally.
@@
  • CEVD (Bose et al. 2023) (CTCE): CEVD modifies the Value Decomposition (VD) paradigm (Sunehag et al. 2018) from CTDE to CTCE. Specifically, it combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, similar to the motivation behind BMG-Q.
For fairness, we modify the state space and reward function to align closely with those used in this paper. Additionally, we also choose a Greedy method as a baseline, which directly matches the order to the closest driver. The training process and testing results are shown in Fig. 2 and Table 1, respectively. In Fig. 2, we illustrate the cumulative reward alongside the number of orders served, as well as the average delivery time, detour time, pickup time, and confirmation time for each order. We observe that the GRPO and OSPO outperform the others primarily by achieving lower pickup and confirmation times, allowing them to serve more orders. In Table 1, it is noteworthy that the OSPO, which has the smallest network size and lowest GPU utilization, achieves the best performance across three testing scenarios. However, we also find that the GRPO and OSPO do not perform as well when the order amount increases to 3,910, which is the only scenario where the order amount exceeds 3,726 in the training set. Notably, the two methods that perform relatively well in this scenario are BMG-Q (Hu, Feng, and Li 2025) and CEVD (Bose et al. 2023). The primary distinction of these methods compared to others is their use of neighbor information when calculating Q-values. In Fig. 2, we also observe that they achieve lower detour times, suggesting better cooperation among agents through neighbor information. These results imply that incorporating more global or neighbor information may be more beneficial in on-pick scenarios. Moving forward, it would be advantageous to explore the model’s performance when the training set includes multiple order amount scenarios.
Ablation Study
@@
  • Deviation Punishment: We examine the impact of including deviation punishment ($\Delta\delta_{t}$ in Eq. 12) to determine whether it helps to maintain property (ii) (Eq. 8).
@@
Figure 3: Result of Ablation Study
@@
Additionally, we observe that OSPO demonstrates overall better performance than GRPO. This is because, although the V-values among agents are generally similar, they have been maintained over a longer time period. In our ride-sharing scenario, the number of served orders may differ slightly among agents at each time step, but this variation can be considered negligible over the long term. Since our simulation lasts only 30 minutes, this difference becomes more pronounced when calculating the advantage (i.e., using the group average cumulative reward as a baseline introduces a slight bias). In contrast, OSPO exhibits greater robustness, as it relies solely on one-step rewards and does not encounter this issue.
Discussions
In this section, we analyze the current limitations and future directions of our work.
  • Application Constraints: As mentioned earlier, our method primarily relies on the property that the V-values of different agents remain similar at any time step (Eq. 8). This limits the application scenarios of OSPO to cooperative homogeneous environments. It cannot be used in cooperative scenarios where some agents must sacrifice their own interests to maximize the global reward. However, this condition can mostly be met in independent MARL, as each agent’s learning goal is to maximize its own long-term reward. In the next step, it is worthwhile to explore a wider range of possible application scenarios for OSPO.
  • Independent Assumption: One important reason for proposing OSPO is to reduce the estimation error of Q-values or V-values in independent MARL scenarios. However, the current training approach of our method still follows independent learning. The challenges of poor cooperation in independent MARL are yet to be addressed. Future work could explore combining OSPO with other methods, such as communication mechanisms or more powerful networks, similar to the approach taken by BMG-Q.
  • Fairness: Since OSPO is fundamentally based on the similar V-value property (Eq. 8), and we incorporate a punishment term (Eq. 12) to encourage the model to adhere to this property, it is important to explore whether this approach contributes to fairness. Recently, assignment fairness and discriminatory scenarios have garnered increasing attention in gig systems (Shi et al. 2021).
Conclusion

### v2

@@ 3.2 GRPO Solution
If we ignore the residual term, we obtain $\hat{\text{A}}_{\pi}(s_{t},u_{t})\approx\frac{1}{\sigma}(\text{Q}_{\pi}(s_{t},u_{t})-\text{V}_{\pi}(s_{t}))$. Compared to the advantage in PPO (Eq. 4), there is only one additional factor $\frac{1}{\sigma}$. According to (kurin2022defense), this can be viewed as normalization for the reward function, contributing to reduced deviation and more stable training. As a result, we consider the advantage of GRPO (Eq. 5) to be a valid replacement for the advantage in PPO (Eq. 4).
In MARL, the concept of agents is analogous to the group in GRPO, prompting us to consider whether we can adapt GRPO to the MARL scenario, given its high efficiency and promising performance. The answer is yes, if the agents are completely homogeneous in a cooperative MARL context, such as in our order dispatch task within the ride-sharing system. Under such circumstances, the advantage function can be modified to:
$\hat{\text{A}}_{\pi}^{\prime}(s_{i,t},u_{i,t})=\frac{1}{\sigma_{t:}}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{i,t+\tau},u_{i,t+\tau})-\mu_{t:}\right)\ ,$ (10)
$\mu_{t:}=\mathbb{E}_{i\in\mathcal{I}}\left[\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{i,t+\tau},u_{i,t+\tau})\right]\ .$
$\sigma_{t:}=\sqrt{\mathbb{E}_{i\in\mathcal{I}}\left[(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{i,t+\tau},u_{i,t+\tau})-\mu_{t:})^{2}\right]}\ .$
In a general MARL scenario, we do not have the property that each step’s reward has a similar range, making it infeasible to utilize $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$ to estimate the V-value as shown in Eq. 9. However, based on the homogeneous property outlined in Eq. 2, we can use the expectation of the reward-to-go for all agents, $\mu_{t:}$, as an approximation of the V-value.
Finally, for the KL regularization term in the GRPO objective (Eq. 3), which originally uses a reference LLM policy $\pi_{ref}$, we innovatively propose to use the best-performing policy checkpoint obtained during training. This serves as a stabilizing anchor, preventing the policy from deviating drastically from known good behavior.
3.3 One Step Policy Optimization (OSPO) Solution
Based on the homogeneous property in Eq. 2, we can recalculate the advantage function as follows:
@@ 3.3 One Step Policy Optimization (OSPO) Solution
If we perform a similar operation to GRPO, using the average reward to replace the expectation, applying reward normalization (kurin2022defense), and ignoring the residual term, we arrive at the advantage function:
@@
This indicates that we can use only the one-step reward for policy optimization, which we term One Step Policy Optimization (OSPO). It is important to note that our method fundamentally differs from myopic methods, even though the advantage function incorporates only one-step rewards. The myopic method focuses solely on maximizing the global reward for the current step, which is not the objective of OSPO. In contrast, OSPO optimizes the policy of all agents simultaneously by increasing the probability of favorable actions while reducing the probability of unfavorable ones. The goal of OSPO remains to maximize the target function defined in Eq. 3 (a variant of the long-term accumulative reward in standard RL), but it does so in a one-step format that leverages the homogeneous properties.
@@
$\tilde{\text{A}}_{\pi}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}\ .$ (13)
This approach not only makes $\mu_{t}$ closer to $\mathbb{E}_{j\in\mathcal{I},s_{j,t},u_{j,t}\sim\pi}[\text{r}(s_{j,t},u_{j,t})]$ in Eq. 11, but also further improves computational efficiency, achieving true one-step computation. (For stability, we propose to ignore samples with a reward of 0 when calculating $\mu$ and $\sigma$. This is feasible since those no-assignment samples do not correspond to a real action and therefore cannot be used for training.)
@@
Second, during training, to better align with the property described in Eq. 2, we propose a penalty term $\Delta\delta_{t}$, where $\delta_{t}$ represents the standard deviation of the current cumulative reward for each agent, and $\Delta\delta_{t}=\delta_{t+1}-\delta_{t}$. This penalty is designed to reduce the utility gap (reflecting the V-value) between agents. Consequently, the final advantage function is expressed as:
$\tilde{\text{A}}_{\pi}^{\prime}(s_{i,t},u_{i,t})=\frac{\text{r}(s_{i,t},u_{i,t})-\mu_{t}}{\sigma_{t}}-\alpha\Delta\delta_{t}\ ,$ (14)
@@
where $\alpha$ is a hyper-parameter. Note that when Eq. 2 is satisfied, $\Delta\delta_{t}$ is close to 0. (We also include the term $\alpha\Delta\delta_{t}$ in the advantage function of GRPO to enhance the homogeneous property.) A detailed algorithm process is provided in Appendix A.
Figure 1: The training process consists of four steps: (1) Calculate the matching probability for each vehicle-order pair; (2) Assign orders by maximizing the probability score; (3) Calculate the group advantage; (4) Train the network using policy gradients.
4 Experiment
4.1 Experiment Setup
@@
To validate our method, we conduct a series of experiments using a real-world ride-hailing dataset from Manhattan, New York City (TLCData). The data collected from 19:00 to 19:30 on July 17, 2024, is used as the training set, which includes 3,726 valid orders. We then evaluate the trained model using data from the entire day of July 18, 2024, selecting the first 30 minutes of data from each hour, resulting in a total of 24 independent tests. An expanded generalization experiment is provided in Appendix B.3, using the data from Queens, New York.
@@
For the simulation, we set the total number of vehicles to 1,000, with each car having a capacity of 3 and a default speed of 60 km/h. The optimal routing is achieved using the OSRM simulator (luxen-vetter-2011). The experiments are conducted using the PyTorch framework (paszke2019pytorch) on a workstation running Windows 11, equipped with an Intel(R) Core(TM) i7-14700KF processor and an NVIDIA RTX 4080 graphics card. Following the setup of pervious works, we set each episode length to 30 minutes, with each step representing 1 minute, resulting in a running time of approximately 40 to 120 seconds for different methods per episode. We maximize the training episodes at 1,000 and evaluate model performance every 10 episodes.
@@
To further illustrate the efficiency of our method, we choose a simple four-layer MLP as the policy network, with 128 units in each hidden layer and LeakyReLU (maas2013rectifier) as the activation function. Considering the variable action space (since the number of orders changes each time), we set the network input as a single vehicle-order pair, ignoring the relationships between orders. Finally, a softmax layer is used to normalize the output probabilities.
@@
During training, our models utilized approximately 3.8GB (OSPO) and 5.2GB (GRPO) of GPU memory, with a batch size of 256. For exploration, we add gradually decreasing random noise, employing the same Binary Symmetric Channel (BSC) noise as in (hu2025bmg; hoppe2024global; enders2023hybrid; al2019deeppool). For optimization, we use the Adam optimizer (kingma2014adam) with an initial learning rate of $10^{-4}$ and a decay rate of 0.99.
@@
In the following, we first compare our GRPO and OSPO ride-sharing methods with previously popular methods of different types, including DTDE, CTDE, and CTCE. Next, we conduct a series of ablation studies to detail the effects of different modules in GRPO and OSPO, comparing them with similar policy gradient-based methods, including Independent PPO (IPPO) (de2020independent), Multi-Agent PPO (MAPPO) (yu2022surprising), and Independent PG (IPG) (sutton1999policy).
@@
Table 1: Comparison of Different Ride Sharing Methods: Bold entries represent the best results.
Method OSPO GRPO DeepPool BMG-Q HIVES Enders et al. CEVD Hoppe et al. Reward Type Local Local Local Local Global Local Global Mixed RL Algorithm OSPO GRPO IDDQN IDDQN QMIX MASAC VD COMA Network Backbone MLP MLP MLP GAT GRU MLP MLP MLP Model Size 20K 20K 20K 117K 16M 118K 23K 167K GPU Occupation (GB) 3.82 5.17 3.97 4.28 6.01 8.19 21.45 9.78 Reward ($\times 10^{3}$) 11.00 ± 2.63 10.90 ± 2.61 10.41 ± 2.34 10.88 ± 2.35 10.70 ± 2.32 9.77 ± 2.41 10.71 ± 2.36 9.93 ± 2.43 Service Rate 1.00 ± 0.00 1.00 ± 0.00 0.99 ± 0.02 1.00 ± 0.01 0.99 ± 0.02 0.99 ± 0.02 1.00 ± 0.01 0.99 ± 0.02 Delivery Time 12.32 ± 0.71 12.40 ± 0.73 12.20 ± 1.43 10.61 ± 0.70 10.81 ± 0.51 12.65 ± 0.52 11.95 ± 0.79 12.54 ± 0.61 Detour Time 1.26 ± 0.41 1.42 ± 0.52 1.24 ± 0.90 0.59 ± 0.59 0.84 ± 0.79 2.72 ± 0.33 1.12 ± 0.85 2.54 ± 0.43 Pickup Time 6.71 ± 0.64 7.33 ± 0.60 8.67 ± 1.23 7.62 ± 0.23 8.13 ± 0.31 8.05 ± 0.18 8.02 ± 0.88 8.10 ± 0.58 Confirmation Time 0.00 ± 0.00 0.00 ± 0.00 0.01 ± 0.03 0.00 ± 0.02 0.00 ± 0.01 0.06 ± 0.20 0.00 ± 0.01 0.02 ± 0.08
4.2 Experiment Results
@@
To illustrate the efficiency of our methods, we compare them with several representative ride-sharing approaches from various categories:
@@
  • DeepPool (al2019deeppool): DeepPool is one of the earliest MARL-based ride-sharing works, based on IDDQN (IQL; van2016deep), in which each agent trains its Q-network separately. To align it with our experimental scenario, we replace the CNN used in original paper with an MLP similar to our methods (without the softmax).
@@
  • BMG-Q (hu2025bmg): To achieve better cooperation among agents, BMG-Q is also based on IDDQN but utilizes Graph Attention Network (GAT) (velickovic2017graph) to capture the relationships among neighboring agents.
@@
  • HIVES (hao2022hierarchical): Based on QMIX (rashid2020monotonic), HIVES tries addressing the curse of dimensionality problem by proposing a novel hierarchical mixing structure.
@@
  • enders2023hybrid: Based on Multi-Agent Soft Actor Critic (MASAC) (haarnoja2018soft), Enders et al. propose allowing agents to choose whether to accept assigned orders, which helps the system serve more valuable orders instead of treating all orders equally.
@@
  • CEVD (bose2023sustainable): CEVD modifies the Value Decomposition (VD) paradigm (sunehag2018value) from CTDE to CTCE. Specifically, it combines the Q-values of each agent with those of their neighbors to create a new type of Q-value, similar to the motivation behind BMG-Q.
  • hoppe2024global: Building on the work of enders2023hybrid, a method based on COMA (foerster2018counterfactual) is developed that utilizes a mixture of global and local rewards. Specifically, the global reward encourages cooperation, while the local reward enhances training efficiency.
For fairness, we modify the state space and reward function to align closely with those used in this paper.
The training process and testing results are presented in Fig. 2 and Table 1, respectively. We illustrate the cumulative reward alongside the number of orders served, as well as the average delivery time, detour time, pickup time, and confirmation time for each order. Detailed explanations of these metrics can be found in Appendix B.1. According to the results, we observe that GRPO and OSPO outperform the others primarily by achieving lower pickup and confirmation times, which allows them to serve more orders. Notably, OSPO, with the smallest network size and lowest GPU utilization, achieves the best performance across most testing scenarios. Please note that the results in Table 1 represent the weighted average values from 24 different episodes, calculated according to order volume, along with the standard deviation. More details can be found in Appendix B.2.
Figure 2: Training Process: Each method is trained five times, and the curve is smoothed using Exponential Moving Average (EMA) with $\alpha=0.1$. The shaded area represents the standard deviation of each method. (For delivery time and detour time, only completed orders are counted, as these metrics are uncertain for unfinished orders.)
4.3 Ablation Study
@@
  • Deviation Punishment: We examine the impact of including deviation punishment ($\Delta\delta_{t}$ in Eq. 14) to determine whether it helps to maintain homogeneous property (Eq. 2).
@@
Figure 3: Training Process of Ablation Study
@@
Additionally, we observe that OSPO demonstrates overall better performance than GRPO. This advantage arises from the nature of our ride-sharing scenario, where the number of served orders may vary slightly among agents and episodes. For instance, an agent may receive a newly assigned order at the last step of the simulation, resulting in a relatively higher reward for that episode. While this variation is negligible over the long term, it becomes more pronounced within the confines of our 30-minute simulation when calculating the advantage based on the average reward in Eq. 8. In contrast, OSPO exhibits greater robustness, as it relies solely on one-step rewards and does not encounter this issue.
5 Conclusion